# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ II.4 The MAX-3-XORSAT problem and approximation-hard instance construction
Given the severe overhead of fault tolerance Babbush et al. (2021), quantum hardware is expected to exhibit enormous prefactor disadvantages as compared to parallel silicon, particularly when the comparison is made to hardware with equivalent financial value (e.g. millions of USD). Quantum algorithms thus have the most promise when the problem is hard or outright impossible for classical machines. NP-hard constraint satisfaction problems are no exception, so when benchmarking a proposed quantum algorithm it is important to ensure that the problem classes we consider are sufficiently hard for classical machines, and in the present NISQ era, exhibit their exponential difficulty at small enough $N$ that numerical simulations of quantum algorithms can demonstrate meaningful improvements. MAX-3-XORSAT problems are ideal for benchmarking quantum algorithms because their exponential difficulty scaling is obvious at small $N$ for both classical and prior quantum approaches, in contrast to other problems where the asymptotic exponential scaling often does not set in until system sizes that are prohibitively large for simulation.
@@
To prototype our quantum algorithms, we thus consider MAX-3-XORSAT Arora and Barak (2009), a well-studied problem that consists in the Pauli basis of a hypergraph of $N_{C}$ three-body constraint terms, as sketched in figure 1:
@@
A constraint is said to be satisfied if, for a given bitstring, $V_{ijk}Z_{i}Z_{j}Z_{k}=1$, and unsatisfied otherwise. This is called a hypergraph because $V_{ijk}$ has three indices rather than the usual two found in graph theory. Thanks to the linearity of the problem, one can use Gaussian elimination to check if a solution exists that satisfies all the constraints in $O\left(N^{3}\right)$ time, but if the problem is not fully satisfiable, finding the lowest energy state(s) is NP hard. A random state satisfies half the constraints on average and is thus energy zero. Further, it was shown by Håstad Håstad (2001) that if the true ground state satisfies a fraction $\left(1-\epsilon\right)$ of the constraints, then finding any configuration that satisfies more than $\left(1/2+\epsilon\right)$ of them is also NP-hard. The hardest instances are thus those with small but finite $\epsilon$, e.g. almost satisfiable problems, as both finding the true ground state, and even finding an approximate solution, is exponentially difficult. Note that if the problem graph is sparse (e.g. $N_{C}/N$ is on the order of 1) finding approximate solutions can still be easy, since one can randomly select a fraction of the constraints $<cN$ (for some $O\left(1\right)$ constant $c$), and solve that new, much easier problem; solutions to this sub-problem will satisfy half the remaining constraints, on average.
@@
To ensure that we are studying problems that are both hard to solve and hard to approximate for all known methods, we consider a family of instances we call planted partial solution problems (PPSPs). To construct a PPSP, we choose a small unsatisfied fraction $\epsilon$ and pick a random hypergraph of $N_{C}\gg N$ unique triplets; we use $\epsilon=0.1$ in all simulations here. We then pick a random bitstring $G$ and randomly select $\left(1-\epsilon\right)N_{C}$ of the constraints to be satisfied in $G$, by picking the sign of $V_{ijk}$ appropriately, with the rest unsatisfied. If $\epsilon$ is small and $N_{C}/N\gg 1$, $G$ will be the problem ground state with very high probability, as the SAT/UNSAT transition for this problem is at $N_{C}/N\sim 0.92$ Dubois and Mandler (2002) and at densities much higher than the SAT/UNSAT threshold ground states for random graphs satisfy $N_{C}/2+O\left(\sqrt{N_{C}}\right)$ constraints. This property also ensures that $G$ is a unique ground state with high probability at large $N$. When we refer to random hypergraph problems throughout this work, we refer to this construction rule: a random, potentially fairly dense, hypergraph where one can optionally randomly chose an anomalously large fraction of constraints to be satisfied by matching the signs to a randomly chosen ground state bitstring.
@@
Our PPSP construction is necessary because truly approximation hard problems–where the practical polynomial time approximation difficulty approaches the random guessing limit of the complexity class separation–are rare in the space of all possible instances. Sufficiently sparse problems are approximation-easy, and for denser random problems one can always find strings that satisfy $N_{C}/2+O\left(\sqrt{N_{C}}\right)$ constraints in polynomial time Håstad and Venkatesh (2002), with a smaller prefactor in front of the $\sqrt{N_{C}}$ than the prefactor in the average satisfied in the ground state. We formulated our PPSP construction to ensure our algorithm was being benchmarked on instances with a plausible claim to true classical approximation hardness. We note that commonly studied 3-regular problems Bellitti et al. (2021); Kowalsky et al. (2021) do not display strong approximation hardness, as they are sparse, and can be solved efficiently if satisfiable. And intriguingly, we show that, by some metrics, the performance of our novel algorithm progressively improves with increasing $N_{C}/N$ in this regime, at constant $\epsilon$.
II.5 Absence of angle fine tuning in this work
@@
Before presenting our algorithm and main results, we want to make one last point about what we mean by QAOA when talking about quantum algorithms. Specifically, following the original proposal Farhi et al. (2014), most studies of QAOA allow the individual angles governing the magnitudes of the driver and problem Hamiltonians, $H_{D}$ and $H_{P}$, to vary independently at each timestep. For finding ground states with high depth circuits, this technique is analogous to the schedule fine tuning concept that dates back to the adiabatic formulation of Grover’s algorithm Roland and Cerf (2002), and yields a quadratic speedup. And for many problems, these algorithms exhibit concentration, where the set of angles that is optimal for one randomly generated instance is close to optimal for another with high probability Farhi et al. (2022); Anshu and Metger (2023). More general formulations such as ADAPT-QAOA Zhu et al. (2022) can potentially offer more significant speedups, albeit with the challenge of a much larger search space for optimizing control parameters.
@@
We avoid this approach in our work, and instead focus only on simple heuristics where the schedule is determined from smooth functions based on intuitive guessing and a small amount of trial and error. We do this for two reasons. First, as argued in Kapit and Oganesyan (2021), quadratic speedups from schedule fine tuning are extremely fragile and unlikely to be viable at large $N$ except in very narrow circumstances. Second, as the truly approximation hard instances of MAX-3-XORSAT (and we suspect, many other CSPs) are extremal in the space of random problems, it is less obvious that concentration arguments would apply to the cases we consider. So even if, for example, these instances exhibit concentration and some set of angles is near-optimal for a specific $\epsilon$ and $N_{C}/N$ scaling, we do not think it can be easily assumed that those angles would generalize to other extremal parametrizations.
The fundamental challenge of approximation hardness is to find, or rule out the existence of, one or more very deep minima in an exponentially large search space. Consequently, optimizing angles based on average energies returned is not expected to improve algorithm performance. If the algorithm finds a state in the basin of attraction of the deep minimum, we can likely halt as a successful approximation has been found, but if it does not, sets of angles that return lower energies for random, uncorrelated shallower minima are not expected to improve the probability of reaching states near the deep minimum. And again, even if a given subset of extremal problems does exhibit concentration for angle optimization, those angles may not generalize to any other extremal part of parameter space.
@@
We therefore restrict all of our simulations in this work to linearly increasing total runtime with $N$, and smooth, simple functions to control the relative magnitudes of $H_{D}$ and $H_{P}$. Unsurprisingly, the best scaling and prefactor choices differ somewhat for QAOA and variations of spectrally folded optimization, and the results here are not optimal for any specific algorithm variation or PPSP subclass, but rather represent a decent choice, found by intuition and trial and error, for a broad set of parameters.
III Spectrally folded quantum optimization algorithms
@@
The core idea of spectral folding (and related, more general spectral deformations) is to modify how the Hamiltonian is applied to the quantum state through the introduction of a filter function. Specifically, the algorithms we consider solve problems through simulating the time evolution of a quantum state, as $\left|\psi\right>\to e^{-iH\left(t\right)dt}\left|\psi\right>$, with the exponentiated Hamiltonian discretized as a series of layers e.g. $e^{iaH_{D}}e^{ibH_{P}}$. The driver Hamiltonian, and any other additional Hamiltonian terms, are not changed by spectral folding, so we will ignore them for now and focus on the problem Hamiltonian itself. Specifically, we write $\left|\psi\right>$ in the computational basis as $\left|\psi\right>=\sum_{m=0}^{2^{N}-1}c_{m}\left|m\right>$, where $m$ is the decimal integer representation of a given bitstring. Then, for (arbitrary) control angle $\gamma$:
@@
In spectral folding, the phase of each component instead advances proportional to an arbitrary function $f$ of the diagonal $H_{P}$,
@@
This can be accomplished by introducing a register of auxiliary qubits, applying a gate sequence that maps the sum of the constraint terms to a fraction of that register to store $E$, using a second fraction of that register to compute $f\left(E\right)$, applying a sequence of controlled-phase gates to advance the phase by $f\left(E\right)$, and then uncomputing the previous steps to return the register to its initial state. The entire process is sketched in figure 2. Provided $f$ is a relatively simple function, this adds a multiplicative overhead which is polylogarithmic in $N$, since $E\left(m\right)$ is bounded by a polynomial in $N$ and each arithmetic operation takes $O\left(\log{N}\right)$ steps. [^2] We note that for any choice of $f$ more complex than multiplying $E$ by a constant (something that does not require auxiliary qubits to begin with), any spatial locality the graph might have is lost in this step, since $E\left(m\right)$ is a global quantity which we are deforming with $f$.
@@
Figure 2: A schematic for implementing spectrally deformed time evolution $U=\exp(if(H_{P})dt)$. A set of gates maps the value of each of the $N_{C}$ constraint to adding or subtracting 1 to a register of ancilla qubits, using binary signed adder circuits controlled by the value of $V_{ijk}Z_{i}Z_{j}Z_{k}$. A sequence of additional gates computes $f\left(E\right)$ from $E$ (likely using more ancilla qubits), and then a set of local $Z$ rotations is applied to the register storing $f\left(E\right)$ to advance the phase of each component of $\left|\psi\right>$ proportionally. The computation of $f\left(E\right)$ and entangling with “constraint-controlled” gates are then uncomputed, returning the ancillas to their initial state and disentangling them from the $N$ primary qubits over which the problem is defined. The net result of this entire process is to enact the operation in Eq. (4), evolving time under an arbitrary function of the diagonal Hamiltonian (transverse field layers and other operations on the primary qubits are not shown). For relatively simple functions, the net overhead of this entire process (compared to enacting $\exp\left(iH_{P}dt\right)$ directly) is polylogarithmic in $N$.
@@
We define spectral warping as any $f$ which applies a nonlinear rescaling of $E$, such as $f\left(E\right)=cE^{2}$, and spectral folding as a choice of $f$ that mirrors $E$ about a specific value, e.g. $f\left(E\right)=\left|E-E_{t}\right|$. The core idea is sketched in figure 3, and these two methods can of course be composed. Incorporating this operation enormously expands the space of quantum optimization algorithms we can define; in this work we focus on two choices, linear and quadratic spectral folding. Specifically, if we choose our problem normalization via including a multiplicative constant [^3] Formally, this choice assumes that we know the fraction of the $N_{C}$ constraints which are satisfied in the ground state, something that we cannot know in advance of running our quantum optimization algorithm! However, we can simply repeatedly run the algorithm with different normalization choices to guess its value, a prefactor overhead of at most $O\left(N_{C}\right)$. so that the ground state energy is $E_{\mathrm{GS}}=-N$, and let $E_{t}=AN$ for a constant $A$, then we define linear and quadratic spectral folding as
$f_{\mathrm{lin}}\left(E\right)=\frac{\left|E+AN\right|}{A},\;\;\;f_{\mathrm{quad}}\left(E\right)=\frac{\left(E+AN\right)^{2}}{A^{2}N}$ (5)
@@
Here, $A<1$ defines the approximation target. And critically, it is defined using the conventions that random states have energy zero, so returning a state with energy $AE_{\mathrm{GS}}$ approximates, by a factor of $A$, the degree to which the true ground state itself improves on random states. These normalization choices ensure that the energy difference between the new ground states of the folded problem, and random states, is $N$ as in the original renormalized problem. Making this choice simplifies the analysis significantly.

### v3

@@ II.4 The MAX-3-XORSAT problem and approximation-hard instance construction
Given the severe overhead of fault tolerance Babbush et al. (2021), quantum hardware is expected to exhibit enormous prefactor disadvantages as compared to parallel silicon, particularly when the comparison is made to hardware with equivalent financial value (e.g. millions of USD). Quantum algorithms thus have the most promise when the problem is hard or outright impossible for classical machines. NP-hard constraint satisfaction problems are no exception, so when benchmarking a proposed quantum algorithm it is important to ensure that the problem classes we consider are sufficiently hard for classical machines, and in the present NISQ era, exhibit their exponential difficulty at small enough $N$ that numerical simulations of quantum algorithms can demonstrate meaningful improvements. MAX-3-XORSAT Arora and Barak (2009) problems are ideal for benchmarking quantum algorithms because their exponential difficulty scaling is obvious at small $N$ for both classical and prior quantum approaches, in contrast to other problems where the asymptotic exponential scaling often does not set in until system sizes that are prohibitively large for simulation.
@@
The MAX-3-XORSAT problem consists of a hypergraph of $N_{C}$ three-body constraint terms, as sketched in figure 1:
@@
A constraint is said to be satisfied if, for a given bitstring, $V_{ijk}Z_{i}Z_{j}Z_{k}=1$, and unsatisfied otherwise. A random state satisfies half the constraints on average and is thus energy zero. This is called a hypergraph because $V_{ijk}$ has three indices rather than the usual two found in graph theory. Thanks to the linearity of the problem, one can use Gaussian elimination to check if a solution exists that satisfies all the constraints in $O\left(N^{3}\right)$ time, but if the problem is not fully satisfiable, finding the lowest energy state(s) is NP hard.
Further, it was shown by Håstad Håstad (2001) that if the true ground state satisfies a fraction $\left(1-\epsilon\right)$ of the constraints, then finding any configuration that satisfies more than $\left(1/2+\epsilon\right)$ of them is also NP-hard (see also Khot and Naor (2007); Allen et al. (2015); d’Orsi and Trevisan (2022)). The hardest instances are thus those with small but finite $\epsilon$, e.g. almost satisfiable problems, as both finding the true ground state, and even finding an approximate solution, is exponentially difficult. Note that if the problem graph is sparse (e.g. $N_{C}/N$ is on the order of 1) finding approximate solutions can still be easy, since one can randomly select a fraction of the constraints $<cN$ (for some $O\left(1\right)$ constant $c$), and solve that new, much easier problem; solutions to this sub-problem will satisfy half the remaining constraints, on average.
@@
To ensure that we are studying problems that are both hard to solve and hard to approximate for all known methods, inspired by Refs. Alekhnovich (2003); Khot and Naor (2007); Allen et al. (2015); Barak et al. (2015); d’Orsi and Trevisan (2022) we consider a family of instances we call planted partial solution problems (PPSPs). To construct a PPSP, we choose a small unsatisfied fraction $\epsilon$ and pick a random hypergraph of $N_{C}\gg N$ unique triplets; we use $\epsilon=0.1$ in all simulations here. We then pick a random bitstring $G$ and randomly select $\left(1-\epsilon\right)N_{C}$ of the constraints to be satisfied in $G$, by picking the sign of $V_{ijk}$ appropriately, with the rest unsatisfied. If $\epsilon$ is small and $N_{C}/N\gg 1$, $G$ will be the problem ground state with very high probability, as the SAT/UNSAT transition for this problem is at $N_{C}/N\sim 0.92$ Dubois and Mandler (2002) and at densities much higher than the SAT/UNSAT threshold ground states for random graphs satisfy $N_{C}/2+O\left(\sqrt{N\times N_{C}}\right)$ constraints. When we refer to random hypergraph problems throughout this work, we refer to this construction rule: a random, potentially fairly dense, hypergraph where one can optionally randomly chose an anomalously large fraction of constraints to be satisfied by matching the signs to a randomly chosen ground state bitstring. We note that d’Orsi and Trevisan (2022) recently discovered a hardness threshold at $N_{C}\propto N^{3/2}$, where at higher densities the problem again becomes amenable to classical optimization.
@@
Our PPSP construction is necessary because truly approximation hard problems–where the practical polynomial time approximation difficulty approaches the random guessing limit of the complexity class separation–are rare in the space of all possible instances. Sufficiently sparse problems are approximation-easy, and for denser random problems one can always find strings that satisfy $N_{C}/2+O\left(\sqrt{N\times N_{C}}\right)$ constraints in polynomial time Håstad and Venkatesh (2002); Barak et al. (2015), with a smaller prefactor in front of the $\sqrt{N\times N_{C}}$ than the prefactor in the average satisfied in the ground state. We formulated our PPSP construction to ensure our algorithm was being benchmarked on instances with a plausible claim to true classical approximation hardness. We note that commonly studied 3-regular problems Bellitti et al. (2021); Kowalsky et al. (2021) do not display strong approximation hardness, as they are sparse, and can be solved efficiently if satisfiable. And intriguingly, we show that, by some metrics, the performance of our novel algorithm progressively improves with increasing $N_{C}/N$ in this regime, at constant $\epsilon$.
Finally it is important to note an important phenomenological difference between truly random problems and random approximation-hard instances drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$, for which our PPSPs are merely an efficient construction method. As already mentioned, for truly random problems with large $d_{C}$, the true ground state $G$ satisfies $\frac{N_{C}}{2}+O\left(\sqrt{N_{C}N}\right)$ constraints, and the system exhibits clustering physics in the sense that there are exponentially many well-separated local minima with similar scaling, e.g. close to the energy of $G$. Finding any one cluster is not difficult, so these instances are not approximation-hard in practice, though finding the optimal solution (which may be highly degenerate) is generally hard. In contrast, for random instances with sufficiently large $d_{C}$ and small $\epsilon$, it is overwhelmingly likely that $G$ is unique, and indeed it may not even be possible for a proper clustering phase to emerge near the ground state energy. However, if we look at energies well above $E_{G}$ we find that these problems do exhibit clustering physics and tend to have exponentially many well-separated local minima satisfying the same $\frac{N_{C}}{2}+O\left(\sqrt{N_{C}N}\right)$ as in random problems. We refer to this as the clustering energy throughout this work; the core phenomenological difference between these instances and random problems is thus that one minimum is extensively lower than all the others, and finding it, or configurations close to it in energy, is the goal of our algorithms. That the clustering energy in PPSPs matches that of random problems is consistent with the fact that deciding if a given instance is a truly random problem or a PPSP (in our language) is itself classically hard Alekhnovich (2003); Allen et al. (2015); d’Orsi and Trevisan (2022).
II.5 Parameter choices for algorithm formulation and benchmarking
Before presenting our main theoretical and numerical results, we want to discuss the principles that guide how we choose parameters when formulating and testing our novel algorithms. Our choices of the runtime scaling, and other parameters such normalizing the ground state to $-N$ in the classical problem, and the timestep $dt$, are motivated by the desire to carefully separate fundamental limitations on algorithm performance from incidental ones. Here, we consider fundamental limitations to be the core physics of the problem, principally exponentially small transition rates; these are the limitations that we wrote this paper to explore and which require new theoretical and algorithmic insight to address. Conversely, incidental limitations are more prosaic effects that degrade performance, such as Trotter error from larger timesteps, or the proliferation of local excitations by varying parameters too quickly. These limitations are more generic, and have easier solutions such as smaller timesteps and slower parameter variation, though determining optimal parameter choices (particularly when noise must be taken into account) can still be challenging. However, it is often not straightforward to determine, in practice, why a given algorithm is not performing as expected, particularly when the only useful information available at the end is a set of probabilities of finding various target states.
The parameter choices throughout this work are guided by the goal of having the most easily interpretable results, to check the basic veracity our theories. We did not try to determine the optimal use of quantum resources for a given problem. By far the most important consideration in this vein is the runtime per shot. Obviously, the choice of optimal runtime scaling in any algorithm is often subtle, and our core conjecture–that random hypergraph MAX-3-XORSAT instances can with high probability be efficiently approximated in quantum polynomial time–is somewhat loose on the degree of the polynomial scaling. Our choice of generally linear scaling is motivated by the fundamental physics of the problems under consideration.
Specifically, all of the algorithms in this work consist of parameter sweeps, where states in a target manifold $\mathcal{T}$ are found through collective phase transitions from a trivial initial state as it energetically crosses the states in $\mathcal{T}$. In the case of multi-stage filtered optimization (MSFO), $\mathcal{T}$ is simply the classical problem ground state, where for the folded optimization algorithms $\mathcal{T}$ includes all states near the approximation target $AE_{GS}$ (both algorithms are detailed in the next section). Following the arguments in Kapit and Oganesyan (2021), the total success probability for runtime $t_{F}\left(N\right)$ is expected to scale as
$P\left(t_{F}\right)\propto\sum_{j\in\mathcal{T}}\frac{\left|\Omega_{0j}^{2}\right|}{W_{j}\left(N\right)}t_{F}\left(N\right),$ (2)
where $\Omega_{0j}$ is the matrix element to transition to state $j$ in $\mathcal{T}$. If the algorithm proceeds by varying dimensionless control parameter $s$, then the energy scale $W_{j}\left(N\right)$ is proportional to the slope of $\partial\left(E_{0}-E_{j}\right)/\partial s$ in the vicinity of the transition
[…]
$\Delta\left(s\right)\propto 2\sqrt{\Omega_{0j}^{2}+W_{j}^{2}\left(s-s_{c}\right)^{2}},$ (4)
and can be extracted by fitting to this form. For the problems and energy scales chosen in this work, all $W_{j}$ are asymptotically $O\left(N\right)$ in all studied cases.
This scaling, in our estimation, immediately sets a floor of $t_{F}\left(N\right)\propto N$. Even in a very “easy” problem where $\Omega_{0j}$ is constant with $N$, for algorithms of this type for any runtime $t_{F}\left(N\right)$ that scales sublinearly $P\left(t_{F}\right)$ will subsequently decay with system size. Further, for sufficiently fast ramps, one needs to consider local diabatic heating that arises from varying Hamiltonian parameters too quickly, as a process where the system successfully transitions into a target state in $\mathcal{T}$ but is then excited out of it by creating local excitations similarly reduces $P\left(t_{F}\right)$, causing it to decay more quickly than Eq. 2 predicts. Fortunately, for a system which is locally gapped on either side of the phase transition (as a PPSP is, though not necessarily with spectral filtering added), linear runtime is generally enough to make this issue negligible.
Runtimes longer than linear, in contrast, present their own interpretational challenges. For the MSFO defined in the next section, the prediction of constant $P\left(t_{F}\right)$ with $t_{F}\left(N\right)\propto N^{1.46}$ comes from a calculation capable of predicting that precise scaling, but for folded optimization with an approximation target $A<1$ (see below for definitions) our predictions are not able to predict the degree of polynomial scaling directly. For instance, the choice of $A\geq 0.6$ in the spectral folding TMA formulation comes from a calculation with many approximations expected to underestimate performance, and the empirical decay exponents for finding approximate solutions below the achievable approximation ratio $q_{a}$ are often quite small. Given that these algorithms are not amenable to more efficient classical simulation techniques such as tensor network methods, owing to their relatively high circuit depths and nonlocal structure, and the need to average over many instances to get good statistics, we were limited to system sizes in the high twenties at most in numerical simulations. This can make it tricky to resolve small exponents from transient polynomial scaling when trying to determine $q_{a}$. Our choice of linear scaling in this respect is further motivated by the fact that in the folded regime, both $\left|\Omega_{0j}\right|^{2}\right>_{j}$ and the number of states in $\mathcal{T}$ tend to exhibit simple exponential decay and growth, respectively, and consequently polynomial prefactors in $P\left(t_{F}\right)$ are expected to be minimized for $t_{F}\left(N\right)\propto N$ since $W_{j}\left(N\right)$ scales linearly as well.
Further, though this is not a rigorous comparison, the very simplest classical improvement on random guessing–greedy or quasi-greedy steepest descent–usually takes $O\left(N\right)$ steps to halt, and if restricted to a constant number of cost function evaluations per shot, will asymptotically return an approximation ratio $q_{a}=0$ regardless of $d_{C}$ and $\epsilon$ as $N\to\infty$. This is true even if the problem is convex, and more complex methods generally involve longer per-shot runtimes. And as remarked earlier, sophisticated mathematical analyses have provided similar worst-case performance bounds for QAOA and TAQC at constant depth for $k$-XORSAT problems Basso et al. (2022); Boulebnane and Montanaro (2022); Anshu and Metger (2023); Benchasattabuse et al. (2023), though these restrictions do not apply if we allow the circuit depth to grow with system size. For all these reasons, we thus used exclusively linear runtimes for all numerical experiments directly probing approximation hardness, and subquadratic runtimes for MSFO as guided by a rigorous theoretical prediction.
@@ II.5 Parameter choices for algorithm formulation and benchmarking
Finally, we want to address the guiding principles we used for tuning the schedules and other hyperparameters of our algorithms. It has been known since the adiabatic formulation of Grover’s algorithm Roland and Cerf (2002) that schedule fine tuning can produce a quadratic speedup over a uniform sweep, and more general variational approaches such as ADAPT-QAOA Zhu et al. (2022) (which use large fine-tuned angles and draw operators at each step from a much larger set) can potentially offer more significant speedups, albeit with the challenge of a much larger search space for optimizing control parameters. And for many problems, these algorithms exhibit concentration, where the set of angles that is optimal for one randomly generated instance is close to optimal for another with high probability Farhi et al. (2022); Anshu and Metger (2023).
@@
We do not employ such fine tuning in this work for four reasons. First, as argued in Kapit and Oganesyan (2021), quadratic speedups from schedule fine tuning are fragile (at least, for gaps that decay exponentially) and unlikely to be viable at large $N$ except in very narrow circumstances. Second, as the truly approximation hard instances of MAX-3-XORSAT (and we suspect, many other CSPs) are extremal in the space of random problems, it is less obvious that concentration arguments would apply to the cases we consider. So even if, for example, these instances exhibit concentration and some set of angles is near-optimal for a specific $\epsilon$ and $d_{C}=N_{C}/N$ scaling, we do not think it can be easily assumed that those angles would generalize to other extremal parametrizations. Third, we are interested in results we can explain, and extrapolate to larger $N$, through analytical arguments, and beyond-quadratic performance improvements from angle fine tuning are very difficult to understand in this respect. And finally, the overhead of fine tuning for specific problem instances or classes can be both severe and difficult to calculate a priori. If such fine tuning is necessary for good algorithmic performance its overhead needs to be carefully accounted for when estimating the degree of quantum speedup. We are not aware of any results showing significant quantum speedups for MAX-3-XORSAT using QAOA and angle fine tuning.
@@
We thus restrict all of our simulations in this work to employ schedules governed by smooth, simple functions to control the relative magnitudes of $H_{D}$ and $H_{P}$. Unsurprisingly, the best scaling and prefactor choices differ somewhat for Trotterized AQC (TAQC) and variations of spectrally filtered optimization, and the results here are not optimal for any specific algorithm variation or PPSP subclass, but rather represent a decent choice, found by intuition and trial and error, for a broad set of parameters.
III Spectrally filtered quantum optimization algorithms
III.1 Key preliminaries – filter functions
@@
The core idea of spectrally filtered quantum optimization is to modify how the Hamiltonian is applied to the quantum state through the introduction of a filter function, somewhat analogous to the methods used in Hastings (2018); Dalzell et al. (2023), though our goals and filtering choices are very different. Specifically, the algorithms we consider solve problems through simulating the time evolution of a quantum state, as $\left|\psi\right>\to e^{-iH\left(t\right)dt}\left|\psi\right>$, with the exponentiated Hamiltonian discretized as a series of layers e.g. $e^{iaH_{D}}e^{ibH_{P}}$. The driver Hamiltonian, and any other additional Hamiltonian terms, are not changed by spectral filtering, so we will ignore them for now and focus on the problem Hamiltonian itself. Specifically, we write $\left|\psi\right>$ in the computational basis as $\left|\psi\right>=\sum_{m=0}^{2^{N}-1}c_{m}\left|m\right>$, where $m$ is the decimal integer representation of a given bitstring. Then, for (arbitrary) control angle $\gamma$, the wavefunction expressed in the computational $Z$ basis evolves as
@@
In spectrally filtered optimization, the phase of each component instead advances proportional to an arbitrary function $f$ of the diagonal $H_{P}$,
@@
This can be accomplished by introducing a register of auxiliary qubits, applying a gate sequence that maps the sum of the constraint terms to a fraction of that register to store $E$, using a second fraction of that register to compute $f\left(E\right)$, applying a sequence of controlled-phase gates to advance the phase by $f\left(E\right)$, and then uncomputing the previous steps to return the register to its initial state. The entire process is sketched in Fig. 2. Provided $f$ is a relatively simple function, this adds a multiplicative overhead which is polylogarithmic in $N$, since $E\left(m\right)$ is bounded by a polynomial in $N$ and each arithmetic operation takes $O\left(\log{N}\right)$ steps. [^3] We note that for any choice of $f$ more complex than multiplying $E$ by a constant (something that does not require auxiliary qubits to begin with), any spatial locality the graph might have is lost in this step, since $E\left(m\right)$ is a global quantity which we are deforming with $f$.
@@
Figure 2: A schematic for implementing spectrally deformed time evolution $U=\exp(if(H_{P})dt)$. A set of gates maps the value of each of the $N_{C}$ constraints to adding or subtracting 1 to a register of ancilla qubits, using binary signed adder circuits controlled by the value of $V_{ijk}Z_{i}Z_{j}Z_{k}$. A sequence of additional gates computes $f\left(E\right)$ from $E$ (likely using more ancilla qubits), and then a set of local $Z$ rotations is applied to the register storing $f\left(E\right)$ to advance the phase of each component of $\left|\psi\right>$ proportionally. The computation of $f\left(E\right)$ and entangling with “constraint-controlled” gates are then uncomputed, returning the ancillas to their initial state and disentangling them from the $N$ primary qubits over which the problem is defined. The net result of this entire process is to enact the operation in Eq. 7, evolving time under an arbitrary function of the diagonal Hamiltonian (transverse field layers and other operations on the primary qubits are not shown). For relatively simple functions, the net overhead of this entire process (compared to enacting $\exp\left(iH_{P}dt\right)$ directly) is polylogarithmic in $N$.
@@
The algorithms we consider in this work use two classes of filter function, which we call nonlinear folding and warping transformations, both of which assume that the classical problem has been normalized  [^4] Formally, this choice assumes that we know the fraction of the $N_{C}$ constraints which are satisfied in the ground state, something that we cannot know in advance of running our quantum optimization algorithm! However, we can simply repeatedly run the algorithm with different normalization choices to guess its value, a prefactor overhead of at most $O\left(N_{C}\right)$. so that $-N\leq E\leq N$. We define the nonlinear folding transformation as:
$F_{f}\left(E,N,A,x\right)\equiv N\,{\rm sign}\left(E\right)\,\left(1-\left|\left|\frac{E}{N}\right|-A\right|^{x}\right).$ (8)
@@
Here, $A\leq 1$ sets the approximation target of this algorithm, and $x$ sets the (continuous) degree of the fold; the choice $F_{f}\left(E,N,1,1\right)$ applied to the normalized $H_{P}$ has no physical or algorithmic consequences. Critically, it is defined using the conventions that random states have energy zero, so returning a state with energy $AE_{\mathrm{GS}}$ approximates, by a factor of $A$, the degree to which the true ground state itself improves on random states. These normalization choices ensure that the energy difference between the new ground states of the folded problem, and random states, is $N$ as in the original renormalized problem. Making this choice simplifies the analysis significantly. We similarly define the warping transformation as:
$F_{w}\left(E,N,w\right)\equiv N^{1-w}\times{\rm sign}\left(E\right)\times\left|E\right|^{w}.$ (9)